A random sample of 388 married couples found that 280 had two or more personality preferences in common. In another random sample of 562 married couples, it was found that only 36 had no preferences in common. Let p1 be the population proportion of all married couples who have two or more personality preferences in common. Let p2 be the population proportion of all married couples who have no personality preferences in common.

(a) Find a 99% confidence interval for p1p2. (Use 3 decimal places.)

lower limit    
upper limit A random sample of 388 married couples found that 280 had two or more personality preferences in common. In another random sample of 562 married couples, it was found that only 36 had no preferences in common. Let p1 be the population proportion of all married couples who have two or more personality preferences in common. Let p2 be the population proportion of all married couples who have no personality preferences in common.

(a) Find a 99% confidence interval for p1p2. (Use 3 decimal places.)

lower limit    
upper limit
Answer #1

p1cap = X1/N1 = 280/388 = 0.7216
p1cap = X2/N2 = 36/562 = 0.0641

Here, , n1 = 388 , n2 = 562
p1cap = 0.7216 , p2cap = 0.0641


Standard Error, sigma(p1cap - p2cap),
SE = sqrt(p1cap * (1-p1cap)/n1 + p2cap * (1-p2cap)/n2)
SE = sqrt(0.7216 * (1-0.7216)/388 + 0.0641*(1-0.0641)/562)
SE = 0.025

For 0.99 CI, z-value = 2.58
Confidence Interval,
CI = (p1cap - p2cap - z*SE, p1cap - p2cap + z*SE)
CI = (0.7216 - 0.0641 - 2.58*0.025, 0.7216 - 0.0641 + 2.58*0.025)
CI = (0.593 , 0.722)